Use the Ksp values to calculate the molar solubility of each of the following compounds in pure water.

Part A

MX ( K sp = 5.30×10−11)

Express your answer in moles per liter.

Part B

Ag2CrO4 (Ksp = 1.12×10−12)

Express your answer in moles per liter.

Part C

Ni(OH)2 (Ksp = 5.48×10−16)

Express your answer in moles per liter.
Answer #1

A)

At equilibrium:
MX        <---->     M+       +          X-     


                     s                  s       


Ksp = [M+][X-]
5.3*10^-11=(s)*(s)
5.3*10^-11= 1(s)^2
s = 7.28*10^-6 M
Answer: 7.28*10^-6 M

B)

At equilibrium:
Ag2CrO4   <---->    2 Ag+     +          CrO42-


                     2s                 s       


Ksp = [Ag+]^2[CrO42-]
1.12*10^-12=(2s)^2*(s)
1.12*10^-12= 4(s)^3
s = 6.542*10^-5 M
Answer: 6.54*10^-5 M

C)

At equilibrium:
Ni(OH)2   <---->     Ni2+     +         2 OH-   


                     s                  2s      


Ksp = [Ni2+][OH-]^2
5.48*10^-16=(s)*(2s)^2
5.48*10^-16= 4(s)^3
s = 5.155*10^-6 M
Answer: 5.16*10^-6 M